All answers must be answered as a formula
1. Calculate the operating cash flow Calculating OCF-Exce HOME INSERTPAGE LAYOUTFORMULAS DATA REVIEWVEW Sign In Paste Clipboard Al T-3A Aligen berEatng Formatting Table Styles E F Chevelle, Inc., has sales of $39,500, costs of $18,400, depreciation expense of $1,900, and interest expense of $1,400. If the tax rate is 35 percent, what is the operating cash flow, or OCF? S 39,500 18.400 1,900 1,400 Sales Depreciation Expense Interest Expense 10 Tax rate 35% 12 13 14 15 16 17 Complete the following analysis. Do not hard code values in your answers. Income Statement Sales Costs Depreciation expense EBIT Interest expense 19 Sheeti Attempt(s)
1. Calculate the operating cash flow. Calculating OCF-Excel ?固 HOME INSERT PAGE LAYOUT FORMULAS DATA REVIEW VIEW Sign In Pasteメ . |田·!ク·A . Alignment: Number. Conditional Format as-Cell Formatting Table Styles B 1 Cells Editing Font A1 A B 10 Interest Expense Tax rate 12 13 35% 15 16 17 18 19 20 21 Complete the following analysis. Do not hard code values in your answers. ncome Statement Sales Costs Depreciation expense EBIT Interest expense EBT 200 24 25 Taxes (35%) Net income Operating cash flow 32 Sheet1 100% Attempt(s)
Answer #1

EBIT=Sales-costs-depreciation expense

=39500-18400-1900

=19200

EBT=EBIT-interest expense

=19200-1400

=17800

Taxes (35%)=EBT*tax rate

=17800*35%

=6230

Net income=EBT-taxes

=17800-6230

=11570

Operating cash flow=EBIT-Taxes+depreciation expense

=19200-6230+1900

=14870

the above is answer..